OpenStudy (republican31):

I need help setting up the problem for Cashews cost $4 per pound and peanuts cost $1 per pound. How many pounds of peanuts should be used with the cashews to get a mixture of 24 pounds that will cost $2 per pound?

OpenStudy (anonymous):

1x + 4(24-x) = 2(24) x + 96 - 4x = 48 x - 4x = 48 - 96 -3x = -48 x = 48/3 x = 16 lb

OpenStudy (anonymous):

Or x + 4(24-x) = 48 (cost of mix) 3x = 48 -> x = 16 ?

OpenStudy (republican31):

Thank you! Sorry I actually ended up figuring it out on my own forgetting that I posted this here. It went the same way as your first answer! Thank you!